# Electron-Phonon interaction and lattice thermal conductivity from metals to 2D Dirac crystals: a review

### v1

@@ 5.3 Effect of first–order electron–phonon interactions on the lattice thermal conductivity
In lightly doped or gated graphene ($n\!\lesssim\!10^{12}\,\mathrm{cm^{-2}}$), the static screening wave vector grows as $q_{s}\!\propto\!k_{F}\!\propto\!\sqrt{n}$ and is therefore small.[91, 92] In this regime the gauge-field (shear) coupling that governs long-wavelength LA/TA scattering is essentially unscreened, while scalar (deformation-potential–like) terms are only weakly screened.[99, 100] If mirror symmetry is broken (by a substrate, gating, ripples, or edges), a linear one-phonon coupling to flexural (ZA) modes is activated with small-$q$ amplitude $\propto q$;[93, 102] under those conditions the ZA-induced electronic scattering can become comparable to the LA/TA channel at room temperature for low $n$. First-principles electron BTE calculations (beyond RTA) find that allowing symmetry-broken, one-phonon ZA processes reduces the room-temperature mobility by on the order of tens of percent in otherwise clean devices, with a weak $n$-dependence for $n\!\lesssim\!10^{12}\,\mathrm{cm^{-2}}$.[99, 102] The weak density trend reflects two facts: (i) at small $k_{F}$ the static screening of scalar terms is modest, and (ii) the dominant low-$q$ matrix elements scale only linearly with $q$, so the overall phase-space weighting does not change dramatically with $n$ in this limit.[92, 99]
At higher doping ($n\!\gtrsim\!10^{13}\,\mathrm{cm^{-2}}$), $q_{s}$ exceeds typical thermal phonon wave vectors ($q\!\sim\!k_{B}T/\hbar v_{s}$ for LA/TA) and strongly suppresses scalar (deformation-potential–like) contributions—including the symmetry-allowed ZA coupling—except in the narrow backscattering window near $q\!\approx\!2k_{F}$.[91, 92] The in-plane acoustic channel, dominated by the unscreened gauge-field coupling, therefore regains clear dominance.[99, 100] In this degenerate regime the intrinsic acoustic-phonon-limited resistivity remains approximately linear in $T$ and only weakly dependent on $n$; consequently, the conductivity $\sigma$ grows roughly $\propto n$ (at fixed $T$), while the mobility $\mu=\sigma/(ne)$ is nearly $n$-independent.[92, 100] By Wiedemann–Franz, the electronic thermal conductivity then scales as $\kappa_{\mathrm{e}}\!\approx\!L_{0}T\,\sigma$ and increases with $n$ through $\sigma$, whereas flexural-phonon contributions to electronic scattering become a negligible correction in clean, highly doped samples.[92]
@@ 5.5 Size, edge roughness, and specularity
with $P$ the specularity parameter (P=1P\!=\!1 mirror-like, P=0P\!=\!0 fully diffuse) and $v_{\nu}$ the branch group velocity.[107] This has been shown in Fig. 6(b). Because $v_{\mathrm{LA}}$ and $v_{\mathrm{TA}}$ exceed the small-$q$ flexural velocity $v_{\mathrm{ZA}}(q)=\partial\omega_{\mathrm{ZA}}/\partial q\propto q$ by an order of magnitude, the same edge roughness (1P)(1-P) imposes a larger additional rate on the in-plane modes.
@@
Overall, rough or narrow geometries preferentially attenuate the fast in-plane LA/TA carriers far more than the slower but densely populated ZA modes, making the ZA branch an efficient thermal channel in graphene nanoribbons over a wide range of sizes and edge specularities.[13, 107, 103]
@@ 5.6 The role of higher order electron-phonon interactions
In the four–particle channel $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$, crystal momentum is conserved modulo a reciprocal–lattice vector $\mathbf{G}$ (Umklapp). Whenever the created phonon lies outside the first Brillouin zone its wave vector folds back as $\mathbf{q}^{\prime\prime}-\mathbf{G}$. A simple enumeration of neighboring zones shows that roughly one half of the nominally allowed $\mathbf{q}^{\prime\prime}$ fall into Umklapp–active sectors where $|\mathbf{q}^{\prime\prime}-\mathbf{G}|\!\ll\!|\mathbf{q}^{\prime\prime}|$; this is depicted in Fig. 7(b). These events divert electronic momentum into lattice recoil and effectively remove about half of the kinematically available outgoing–phonon phase space for $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$, thereby lowering the corresponding four–particle rate and imposing an intrinsic cap on the associated phonon mean–free path [35]. Taken together, (i) the need for a Lindhard (not TF) dielectric with a $2k_{F}$ cusp for acoustic phonons, and (ii) the rising importance of $\mathrm{E}\!\to\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$ and $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$ when $E_{\mathrm{F}}\!\sim\!k_{B}T$, are derived and quantified specifically for 2D Dirac crystals.
@@
and the e–ph–limited transport is governed by the in–plane acoustic branches (LA/TA), consistent with the classic in–plane graphite treatment [93, 108, 109]. In case (ii) (GNRs), size and edge roughness renormalize branch contributions: diffusive edges and finite length suppress LA/TA more strongly than ZA, so the relative ZA share of $\kappa_{\mathrm{ph}}$ can increase as ribbons narrow or edges roughen (see Fig. 6(a)); nonetheless, for long, high–quality ribbons with nearly specular edges the LA/TA modes remain the dominant heat carriers [107, 93, 98, 13, 103, 110]. Accordingly, the four–particle channel $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$ that we formulate below is written explicitly for the LA/TA branches. Flexural (ZA/ZO) contributions are excluded by symmetry in $\sigma_{h}$–preserving monolayers, and are subdominant under the ribbon conditions we emphasize here.
Figure 7: (a) The four-particle e-ph interaction consisting of the annihilation of a pair of electrons and phonons and the creation of a new pair, $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$, with a Fermion propagator in the middle. (b) The wave vector of the outgoing phonon scattered to the 1st, 4th, 5th, and 8th neighboring lattice sites remain approximately the same while the wave vector of the phonon scattered to the 2nd, 3rd, 6th, and 7th neighboring lattice sites will be much smaller than the annihilated phonon since its wave vector will be hugely affected by the Umklapp process rendering the process inelastic. Figure taken from Kazemian et al.[35].
@@
where is the speed of sound and $\theta$ is the angle of collision between the annihilated phonon and electron and $\chi(q)$ is the dynamic Lindhard dielectric response function.[27] Knowing the e-ph lifetime $\tau_{(e-ph)}$ one can easily calculate the e-ph thermal conductivity
@@
Figure 8: (a,b) The e-ph thermal conductivity with respect to the Fermi energy of a 2D Dirac crystal with $\alpha=10^{-7}$. The variable $\alpha$ is dependent on the unit cell area, mass, and the ion charge of the 2D Dirac crystal. For a 2D Dirac crystal such as graphene with a unit cell area of $\Omega\approx 3.2*10^{-16}$ we have $\alpha\approx 10^{-7}$. We observe that when we incorporate the four-particle process in the calculation of the e-ph thermal conductivity, the rate of increase in slope is significantly reduced at lower Fermi energies compared to when only the three-particle process is considered. This finding emphasizes the significance of higher-order e-ph interactions in the investigation of 2D Dirac crystals, particularly at low Fermi energies. Figure taken from Kazemian et al.[35].
@@ 5.7 Summary and open challenges
Carrier density, lattice symmetry, strain, and finite geometry jointly dictate how e-ph interactions limit transport in 2D Dirac crystals. At low carrier densities ($n\!\lesssim\!10^{12}\,\mathrm{cm}^{-2}$), screening is weak ($q_{s}\!\propto\!k_{F}$), linear one-phonon flexural (ZA) coupling is activated whenever mirror symmetry is broken (by substrates, curvature, or edges), and ZA-induced electronic scattering can become comparable to the in-plane LA/TA channel, reducing intrinsic mobility by on the order of tens of percent in otherwise clean devices.[93, 99, 102] At higher densities ($n\!\gtrsim\!10^{13}\,\mathrm{cm}^{-2}$), Thomas-Fermi/Lindhard screening suppresses scalar (deformation-potential-like) terms and restricts backscattering to a narrow window near $q\!\approx\!2k_{F}$, restoring in–plane dominance. In this degenerate regime In degenerate, acoustic-phonon-limited graphene the resistivity is approximately linear in $T$ and only weakly dependent on $n$; hence at fixed $T$ the conductivity is roughly $n$-independent (and the mobility scales as $\mu\!\propto\!1/n$), unless additional mechanisms (e.g. screening of a scalar deformation potential or multi-channel scattering) introduce an $n$-dependence.[92, 100, 99] Uniform tensile strain introduces an in-plane tension $\sigma=Y\epsilon$ that linearizes the ZA dispersion below a crossover $q_{\ast}=\sqrt{\sigma/\kappa_{b}}$, converting the constant (quadratic) ZA density of states into $D(\omega)\!\propto\!\omega$, raising the long–wavelength ZA group velocity, lengthening ZA lifetimes, and increasing the intrinsic lattice thermal conductivity of suspended membranes (with the gain ultimately limited by sample size and boundaries).[98, 13, 103] If horizontal–mirror symmetry ($\sigma_{h}$) is exact and all atoms lie in the mirror plane (e.g. free-standing graphene), the first-order ZA/ZO e-ph matrix elements vanish. Once symmetry is broken by edges or substrates, linear ZA coupling reappears. In graphene nanoribbons, boundary scattering modeled via Ziman/Casimir specularity suppresses LA/TA strongly and can raise the relative ZA share as ribbons shorten or edges roughen, although in long, high-quality ribbons the in-plane branches typically remain the dominant heat carriers.[98, 107, 108] Furthermore, it has been shown that in 2D Dirac crystals, at elevated temperatures and low Fermi energies, the E $\rightarrow$ EP process has the potential to partially or entirely nullify the EP $\rightarrow$ E process. It therefore becomes imperative to investigate higher order e-ph interactions in order to comprehend the overall dynamics accurately. Findings demonstrate that the impact of the four-particle process, $\leftrightarrow EP$, on both phonon scattering rate and e-ph thermal conductivity. Umklapp further prunes the available phase space in this four-particle channel, modifying phonon lifetimes and the e-ph contribution to $\kappa$.[35]
@@
Looking forward, several challenges remain within Dirac crystals. A first priority is to extend coupled e-ph transport formalisms to time-dependent drives in the ultrafast regime of hot carriers and nonthermal phonons, while enforcing Kelvin-Onsager reciprocity. For graphene this requires marrying first-principles e-ph matrix elements with nonequilibrium carrier dynamics and the symmetry constraints specific to ZA/ZO modes.[48, 91, 108] Second, beyond-RTA, fully iterative/variational solutions of the Peierls-Boltzmann equation with mode-to-mode couplings and higher order scattering are needed to quantify drag and boundary effects as device dimensions approach intrinsic mean–free paths.[13] Third, accurate e-ph workflows for correlated or magnetic Dirac crystals will require beyond-DFT electronic structures (hybrid functionals, $GW$, or DMFT) so that EPC rests on realistic quasiparticles.[48, 91] Fourth, systematic experimental benchmarks on graphene and related Dirac membranes TDTR and micro-Raman on ribbons and strained, suspended stacks with known edge specularity, curvature, substrate dielectric environment, and carrier density are still scarce and essential to validate higher order and coupled BTE predictions.[98] Finally, device scale multiphysics that embeds first principles e-ph lifetimes (including ZA symmetry rules and higher-order channels) into electro-thermal solvers is needed to translate these microscopic insights into quantitative design rules for Dirac transistors, interconnects, and thermoelectric elements.[35, 98]

### v2

@@ 5.4 Effect of first–order electron–phonon interactions on the lattice thermal conductivity
In lightly doped or gated graphene ($n\!\lesssim\!10^{12}\,\mathrm{cm^{-2}}$), the static screening wave vector grows as $q_{s}\!\propto\!k_{F}\!\propto\!\sqrt{n}$ and is therefore small.[CastroNeto2009_RMP, DasSarma2011_RMP] In this regime the gauge-field (shear) coupling that governs long-wavelength LA/TA scattering is essentially unscreened, while scalar (deformation-potential–like) terms are only weakly screened.[Sohier2014_PRB, Park2014_NanoLett] If mirror symmetry is broken (by a substrate, gating, ripples, or edges), a linear one-phonon coupling to flexural (ZA) modes is activated with small-$q$ amplitude $\propto q$;[Mariani2008_PRL, Gunst2017_PRL] under those conditions the ZA-induced electronic scattering can become comparable to the LA/TA channel at room temperature for low $n$. First-principles electron BTE calculations (beyond RTA) find that allowing symmetry-broken, one-phonon ZA processes reduces the room-temperature mobility by on the order of tens of percent in otherwise clean devices, with a weak $n$-dependence for $n\!\lesssim\!10^{12}\,\mathrm{cm^{-2}}$.[Sohier2014_PRB, Gunst2017_PRL] The weak density trend reflects two facts: (i) at small $k_{F}$ the static screening of scalar terms is modest, and (ii) the dominant low-$q$ matrix elements scale only linearly with $q$, so the overall phase-space weighting does not change dramatically with $n$ in this limit.[DasSarma2011_RMP, Sohier2014_PRB]
At higher doping ($n\!\gtrsim\!10^{13}\,\mathrm{cm^{-2}}$), $q_{s}$ exceeds typical thermal phonon wave vectors ($q\!\sim\!k_{B}T/\hbar v_{s}$ for LA/TA) and strongly suppresses scalar (deformation-potential–like) contributions—including the symmetry-allowed ZA coupling—except in the narrow backscattering window near $q\!\approx\!2k_{F}$.[CastroNeto2009_RMP, DasSarma2011_RMP] The in-plane acoustic channel, dominated by the unscreened gauge-field coupling, therefore regains clear dominance.[Sohier2014_PRB, Park2014_NanoLett] In this degenerate regime the intrinsic acoustic-phonon-limited resistivity remains approximately linear in $T$ and only weakly dependent on $n$; consequently, the conductivity $\sigma$ grows roughly $\propto n$ (at fixed $T$), while the mobility $\mu=\sigma/(ne)$ is nearly $n$-independent.[DasSarma2011_RMP, Park2014_NanoLett] By Wiedemann–Franz, the electronic thermal conductivity then scales as $\kappa_{\mathrm{e}}\!\approx\!L_{0}T\,\sigma$ and increases with $n$ through $\sigma$, whereas flexural-phonon contributions to electronic scattering become a negligible correction in clean, highly doped samples.[DasSarma2011_RMP]
@@ 5.6 Size, edge roughness, and specularity
with $P$ the specularity parameter (P=1P\!=\!1 mirror-like, P=0P\!=\!0 fully diffuse) and $v_{\nu}$ the branch group velocity.[Shen2014_JAP] This has been shown in Fig. 7(b). Because $v_{\mathrm{LA}}$ and $v_{\mathrm{TA}}$ exceed the small-$q$ flexural velocity $v_{\mathrm{ZA}}(q)=\partial\omega_{\mathrm{ZA}}/\partial q\propto q$ by an order of magnitude, the same edge roughness (1P)(1-P) imposes a larger additional rate on the in-plane modes.
@@
Overall, rough or narrow geometries preferentially attenuate the fast in-plane LA/TA carriers far more than the slower but densely populated ZA modes, making the ZA branch an efficient thermal channel in graphene nanoribbons over a wide range of sizes and edge specularities.[Lindsay2014_PRB, Shen2014_JAP, Kuang2016_IJHMT]
@@ 5.7 The role of higher order electron-phonon interactions
In the four–particle channel $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$, crystal momentum is conserved modulo a reciprocal–lattice vector $\mathbf{G}$ (Umklapp). Whenever the created phonon lies outside the first Brillouin zone its wave vector folds back as $\mathbf{q}^{\prime\prime}-\mathbf{G}$. A simple enumeration of neighboring zones shows that roughly one half of the nominally allowed $\mathbf{q}^{\prime\prime}$ fall into Umklapp–active sectors where $|\mathbf{q}^{\prime\prime}-\mathbf{G}|\!\ll\!|\mathbf{q}^{\prime\prime}|$; this is depicted in Fig. 8(b). These events divert electronic momentum into lattice recoil and effectively remove about half of the kinematically available outgoing–phonon phase space for $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$, thereby lowering the corresponding four–particle rate and imposing an intrinsic cap on the associated phonon mean–free path [kazemian2024influence]. Taken together, (i) the need for a Lindhard (not TF) dielectric with a $2k_{F}$ cusp for acoustic phonons, and (ii) the rising importance of $\mathrm{E}\!\to\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$ and $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$ when $E_{\mathrm{F}}\!\sim\!k_{B}T$, are derived and quantified specifically for 2D Dirac crystals.
@@
and the e–ph–limited transport is governed by the in–plane acoustic branches (LA/TA), consistent with the classic in–plane graphite treatment [Mariani2008_PRL, Alidoosti2022_sigmah, Klemens2000_GraphiteAplane]. In case (ii) (GNRs), size and edge roughness renormalize branch contributions: diffusive edges and finite length suppress LA/TA more strongly than ZA, so the relative ZA share of $\kappa_{\mathrm{ph}}$ can increase as ribbons narrow or edges roughen (see Fig. 6(a)); nonetheless, for long, high–quality ribbons with nearly specular edges the LA/TA modes remain the dominant heat carriers [Shen2014_JAP, Mariani2008_PRL, Nika2017_RPP, Lindsay2014_PRB, Kuang2016_IJHMT, liu2014anomalous]. Accordingly, the four–particle channel $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$ that we formulate below is written explicitly for the LA/TA branches. Flexural (ZA/ZO) contributions are excluded by symmetry in $\sigma_{h}$–preserving monolayers, and are subdominant under the ribbon conditions we emphasize here.
Figure 8: (a) The four-particle e-ph interaction consisting of the annihilation of a pair of electrons and phonons and the creation of a new pair, $\mathrm{EP}\!\leftrightarrow\!\mathrm{E}^{\ast}\mathrm{P}^{\ast}$, with a Fermion propagator in the middle. (b) The wave vector of the outgoing phonon scattered to the 1st, 4th, 5th, and 8th neighboring lattice sites remain approximately the same while the wave vector of the phonon scattered to the 2nd, 3rd, 6th, and 7th neighboring lattice sites will be much smaller than the annihilated phonon since its wave vector will be hugely affected by the Umklapp process rendering the process inelastic. Figure taken from Kazemian et al.[kazemian2024influence].
@@
where is the speed of sound and $\theta$ is the angle of collision between the annihilated phonon and electron and $\chi(q)$ is the dynamic Lindhard dielectric response function.[kazemian2023dynamic] Knowing the e-ph lifetime $\tau_{(e-ph)}$ one can easily calculate the e-ph thermal conductivity
@@
Figure 9: (a,b) The e-ph thermal conductivity with respect to the Fermi energy of a 2D Dirac crystal with $\alpha=10^{-7}$. The variable $\alpha$ is dependent on the unit cell area, mass, and the ion charge of the 2D Dirac crystal. For a 2D Dirac crystal such as graphene with a unit cell area of $\Omega\approx 3.2*10^{-16}$ we have $\alpha\approx 10^{-7}$. We observe that when we incorporate the four-particle process in the calculation of the e-ph thermal conductivity, the rate of increase in slope is significantly reduced at lower Fermi energies compared to when only the three-particle process is considered. This finding emphasizes the significance of higher-order e-ph interactions in the investigation of 2D Dirac crystals, particularly at low Fermi energies. Figure taken from Kazemian et al.[kazemian2024influence].
5.8 Twisted graphene
Twist-engineered graphene hosts flat moiré minibands with enhanced density of states and modified screening, which can amplify e-ph interactions and open new scattering channels set by moiré reciprocal vectors. Microscopically, several works have shown that phonons can mediate appreciable intervalley attraction in magic-angle twisted bilayer graphene (MATBG), potentially supporting superconductivity with $s$- or $d$-wave character depending on competing interactions[Wu2018_PRL_TBGSC, Lian2019_PRL_TBGSC]. Atomistic calculations further indicate strong e-ph coupling, pronounced electron-hole asymmetry, and nonadiabatic effects in MATBG[Choi2018_PRB_TBG_EPC]. Recent experiments report direct signatures of strong intervalley e-ph coupling in MATBG, consistent with these theoretical expectations[Chen2024_Nature_StrongEPC].
On the lattice side, the moiré superlattice reconstructs acoustic/optical branches into moiré phonon bands with mini-gaps and collective domain-wall vibrations, and similar effects arise in aligned graphene/hexagonal boron nitride (h-BN) heterostructures[Koshino2019_PRB_MoirePhonons, Krisna2023_PRB_GhBN_MoirePhonons]. These modes reshuffle the available phase space for e-ph and ph-ph scattering, enable “moiré-Umklapp” momentum relaxation via small wave vectors, and can couple efficiently to flat-band carriers. In supported devices, remote interfacial phonons from substrates (e.g., h-BN/oxides) may further influence resistivity and hot-phonon dynamics; accounting for these environment-dependent channels is increasingly important in quantitative transport.
From a thermal-transport standpoint, moiré systems sit at the interface between simple crystals and complex unit cells with many near-degenerate modes. In regimes where phonon linewidths become comparable to interbranch splittings, coherences/tunneling contributions to $\kappa$ (beyond Peierls populations) may become relevant. The unified formulation $\kappa=\kappa^{\mathrm{P}}+\kappa^{\mathrm{C}}$ provides the appropriate extension in such cases (see our Sec. 1.2.2 note) [Simoncelli2019_NatPhys].
@@ 5.9 Summary and open challenges
Carrier density, lattice symmetry, strain, and finite geometry jointly dictate how e-ph interactions limit transport in 2D Dirac crystals. At low carrier densities ($n\!\lesssim\!10^{12}\,\mathrm{cm}^{-2}$), screening is weak ($q_{s}\!\propto\!k_{F}$), linear one-phonon flexural (ZA) coupling is activated whenever mirror symmetry is broken (by substrates, curvature, or edges), and ZA-induced electronic scattering can become comparable to the in-plane LA/TA channel, reducing intrinsic mobility by on the order of tens of percent in otherwise clean devices.[Mariani2008_PRL, Sohier2014_PRB, Gunst2017_PRL] At higher densities ($n\!\gtrsim\!10^{13}\,\mathrm{cm}^{-2}$), Thomas-Fermi/Lindhard screening suppresses scalar (deformation-potential-like) terms and restricts backscattering to a narrow window near $q\!\approx\!2k_{F}$, restoring in–plane dominance. In this degenerate regime In degenerate, acoustic-phonon-limited graphene the resistivity is approximately linear in $T$ and only weakly dependent on $n$; hence at fixed $T$ the conductivity is roughly $n$-independent (and the mobility scales as $\mu\!\propto\!1/n$), unless additional mechanisms (e.g. screening of a scalar deformation potential or multi-channel scattering) introduce an $n$-dependence.[DasSarma2011_RMP, Park2014_NanoLett, Sohier2014_PRB] Uniform tensile strain introduces an in-plane tension $\sigma=Y\epsilon$ that linearizes the ZA dispersion below a crossover $q_{\ast}=\sqrt{\sigma/\kappa_{b}}$, converting the constant (quadratic) ZA density of states into $D(\omega)\!\propto\!\omega$, raising the long–wavelength ZA group velocity, lengthening ZA lifetimes, and increasing the intrinsic lattice thermal conductivity of suspended membranes (with the gain ultimately limited by sample size and boundaries).[Nika2017_RPP, Lindsay2014_PRB, Kuang2016_IJHMT] If horizontal–mirror symmetry ($\sigma_{h}$) is exact and all atoms lie in the mirror plane (e.g. free-standing graphene), the first-order ZA/ZO e-ph matrix elements vanish. Once symmetry is broken by edges or substrates, linear ZA coupling reappears. In graphene nanoribbons, boundary scattering modeled via Ziman/Casimir specularity suppresses LA/TA strongly and can raise the relative ZA share as ribbons shorten or edges roughen, although in long, high-quality ribbons the in-plane branches typically remain the dominant heat carriers.[Nika2017_RPP, Shen2014_JAP, Alidoosti2022_sigmah] Furthermore, it has been shown that in 2D Dirac crystals, at elevated temperatures and low Fermi energies, the E $\rightarrow$ EP process has the potential to partially or entirely nullify the EP $\rightarrow$ E process. It therefore becomes imperative to investigate higher order e-ph interactions in order to comprehend the overall dynamics accurately. Findings demonstrate that the impact of the four-particle process, $\leftrightarrow EP$, on both phonon scattering rate and e-ph thermal conductivity. Umklapp further prunes the available phase space in this four-particle channel, modifying phonon lifetimes and the e-ph contribution to $\kappa$.[kazemian2024influence]
@@
Looking forward, several challenges remain within Dirac crystals. A first priority is to extend coupled e-ph transport formalisms to time-dependent drives in the ultrafast regime of hot carriers and nonthermal phonons, while enforcing Kelvin-Onsager reciprocity. For graphene this requires marrying first-principles e-ph matrix elements with nonequilibrium carrier dynamics and the symmetry constraints specific to ZA/ZO modes.[Giustino2017_RMP, CastroNeto2009_RMP, Alidoosti2022_sigmah] Second, beyond-RTA, fully iterative/variational solutions of the Peierls-Boltzmann equation with mode-to-mode couplings and higher order scattering are needed to quantify drag and boundary effects as device dimensions approach intrinsic mean–free paths.[Lindsay2014_PRB] Third, accurate e-ph workflows for correlated or magnetic Dirac crystals will require beyond-DFT electronic structures (hybrid functionals, $GW$, or DMFT) so that EPC rests on realistic quasiparticles.[Giustino2017_RMP, CastroNeto2009_RMP] Fourth, systematic experimental benchmarks on graphene and related Dirac membranes TDTR and micro-Raman on ribbons and strained, suspended stacks with known edge specularity, curvature, substrate dielectric environment, and carrier density are still scarce and essential to validate higher order and coupled BTE predictions.[Nika2017_RPP] Fifth, device scale multiphysics that embeds first principles e-ph lifetimes (including ZA symmetry rules and higher-order channels) into electro-thermal solvers is needed to translate these microscopic insights into quantitative design rules for Dirac transistors, interconnects, and thermoelectric elements.[kazemian2024influence, Nika2017_RPP] Finally, for twisted (moiré) graphene, a practical route to quantitative charge and heat transport is to combine DFPT with Wannier–interpolation (e.g., EPW) to obtain e–ph matrix elements on dense meshes in the moiré Brillouin zone, together with moiré phonon calculations that capture band folding, mini–Umklapp processes, and possible remote interfacial phonons from encapsulation/substrates. Transport can then be treated within the Peierls–BTE (population) framework and, where quasi-degenerate branches make coherences relevant, augmented by the unified/coherence correction $\kappa=\kappa^{\mathrm{P}}+\kappa^{\mathrm{C}}$. A fully rigorous microscopic picture in these low-dimensional Dirac systems further benefits from (i) environment-dependent screening (gates/encapsulation), (ii) nonadiabatic effects associated with flat bands, (iii) higher-order anharmonicity (four-phonon) at elevated temperatures, and (iv) a consistency check on Migdal’s approximation when strong e–e correlations coexist with sizable e-ph coupling. Together, these elements connect symmetry breaking, many-body screening, and higher-order scattering to the full hierarchy of transport coefficients and nonequilibrium response functions in twisted/stacked 2D materials.